OpenStudy (anonymous):

Find the equation of the tangent to the graph of y+f^(-1) at x=1 where f(x)= x^(3)+x+1. Can someone show me the steps to this please. Do I first find the inverse function and then get the derivative of that and then plug in x=1. I tried a couple of ways but I keep confusing myself

OpenStudy (asnaseer):

This might help you: http://en.wikipedia.org/wiki/Inverse_functions_and_differentiation

OpenStudy (anonymous):

This is the part that's confusing me. "The graph of: " \[y + f ^{-1}\] What is y? And how is that graphable? That's just an expression. What is the argument for f inverse? Is it f inverse of x? f inverse of y?

OpenStudy (anonymous):

sorry that should say y=f^(-1) not +

OpenStudy (asnaseer):

Did the link help you?

OpenStudy (anonymous):

sort of but I'm still a little messed up on my steps

OpenStudy (anonymous):

Okay that makes more sense. Can I assume it's \[f ^{-1}(x)\ ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The steps you've listed make sense, but finding the explicit inverse function seems pretty tough, eh?

OpenStudy (anonymous):

yea i seem to get mixed up in the numbers and im having a hard time finding the actual inverse

OpenStudy (anonymous):

I don't think that finding the explicit inverse function is actually the solution in this case. It's simple to just look at the specific point give, x=1

OpenStudy (anonymous):

and how do i use that point

OpenStudy (anonymous):

Okay, so the function \[f^{-1}(x)\] takes in outputs of f(x) and puts out the corresponding input of f(x) that gives that output. So: \[f^{-1}(1)\] Will be the x that gives f(x) = 1. By inspection, f(0) = 0^3 + 0 +1 = 1, so \[f^{-1}(1) = 0\]

OpenStudy (anonymous):

F(x) and f inverse look like: